The average accident on a certain road section is at 2.4 per day. What is tge probability that there will be 12 accidents in a week?


1
Expert's answer
2021-05-31T11:34:43-0400

"P(X=x) = \frac{e^{-λ}λ^x}{x!} \\

λ=2.4 \times 7 = 16.8 \\

P(X=12) = \frac{e^{-16.8}16.8^12}{12!} = 0.0533"